perform the indicated calculation
9C4
___
12C4

I don'tunderstand what the question is asking.

Answer by user_dude2008(1862)

9 C 4 = 9!/4!(9-4)! = 9!/(4!5!) = (9*8*7*6*5!)/(4!5!) = (9*8*7*6)/(4*3*2*1) = 126


12 C 4 = 12!/4!(12-4)! = 12!/(4!8!) = (12*11*10*9*8!)/(4!8!) = (12*11*10*9)/(4*3*2*1) = 495



(9 C 4)/(12 C 4) = 126/495 = 14/55 = 0.25454545